You own 18%, or 200,000 shares of Doing Stuff, Inc. Your investment is currently worth $1,500,000. The Company sells 75,000 shares at $6.50/per shares and you do not buy any more shares.

What is the pro forma number of shares (ie total including new shares issued)? (answer in whole numbers. do not round to millions)

What is the pro forma market value of the firm (ie including the new capital raise)? (answer in whole numbers. do not round to millions)

What is the pro forma (new) value of your original investment after this latest round of financing? (answer in whole dollars - do not round to millions)

How much did the value of your original investment change? (calculate as a % but express as number, but really a % - ie if 20%, write as 20)  
Answer #1

1) Total number of shares before new shares are issued

=200,000 shares/18% = 1,111,111 shares

Price of share before issue of new share = 1500,000/200000 = 7.5$ per share

Total number of shares after issue of new share = 1,111,111+75000 = 1,186,111 shares

2) Pro forma market value of firm = number of shares*market price

=(1,111,111*7.5) + (75000*6.5)

=8333,332.5 +487500

=8,820,832.5$

3) pro forma (new) value of your original investment = 200,000 * (8820832.5/1186111)

=200,000*7.44

=1,487,353.63

4) Change in value = 1500,000-1487353.63 / 1500,000

=12646.37/1500,000

=0.84%